There are two consecutive odd integers. When the difference between the larger integer and three is tripled, the result is twenty-five less than five times the smaller integer. Find the integers.
Answer by stanbon(75887)

There are two consecutive odd integers. When the difference between the larger integer and three is tripled, the result is twenty-five less than five times the smaller integer. Find the integers.
----
1st: 2x-1
2nd: 2x+1
----
Equation:
3(2x+1-3) = 5(2x-1)-25
-----
3(2x-2) = 10x-30
6x-6 = 10x-30
4x = 24
x = 6
---
1st: 2x-1 = 11
2nd: 2x+1 = 13
===================
Cheers,
Stan H.
============